OpenStudy (anonymous):

Line m has no y-intercept, and its x-intercept is (3, 0). Line n has no x-intercept, and its y-intercept is (0, -4). The equation of line m is ???? , and the equation of line n is ????

OpenStudy (anonymous):

you need to draw out a graph for this

OpenStudy (anonymous):

i mean a coordinate plane

OpenStudy (anonymous):



OpenStudy (anonymous):

idk what to do w/ the coordinate plane though

OpenStudy (anonymous):

the question is asking for a y=mx+b equation right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well then there is none because horizontal and vertical lines have an undefined slope

OpenStudy (anonymous):

tru

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then maybe its not asking y=mx+b

OpenStudy (mrnood):

if it has no y intercept then it is a vertical line and has the equation x=k (k is constant) if it has no x intercpt then it is a horizontal line and has the equation y=k (k is constant) you can get the values of the constant from the points the line passes through

OpenStudy (anonymous):

@.Gjallarhorn. its x=3 bc x will always be 3 and its y=-4 because y will always =-4

OpenStudy (anonymous):

@mathsims read above

OpenStudy (mrnood):

or you can let someone else give oyu the answer @DaBest21 it is against the code of conduct to give direct answers

OpenStudy (anonymous):

i explained

OpenStudy (anonymous):

He explained. Thanks a bunch to everyone who did:)

OpenStudy (anonymous):

Don"t forget to medal and fan if you already haven't. Plus mention me for more help @mathsims